# Dream to Fly: Model-Based Reinforcement Learning for Vision-Based Drone Flight

### v1

@@ III-E Model-based Reinforcement Learning: DreamerV3
III-E1 World Model training
@@
  • Recurrent Sequence Model: A recurrent sequence model, parameterized by the recurrent state $\bm{h}_{k}$, predicts the evolution of the latent representation. Given the previous latent state $[h_{k-1},z_{k-1}]$ and the action $\bm{a}_{k-1}$, it predicts the current recurrent state $\bm{h}_{k}$.
@@
III-E2 Actor-Critic Training
@@
Figure 3: Reward evolution by number of steps for three different tracks: Circle track, Kidney Track and Figure 8 track. The training performance of DreamerV3 is shown in blue, and for PPO in orange. We show that while PPO is not able to achieve any considerable training in 10 million environment interactions, DreamerV3 is able to train to convergence for the three tracks.
III-F Quadrotor Dynamics
In this section, we describe the nominal dynamics $\vec{f}(\vec{x},\vec{u})$, where $\vec{x}\in\mathbb{R}^{10}$ is the state of the quadrotor and $\vec{u}\in\mathbb{R}^{4}$ is the input to the system. The state of the quadrotor is given by $\vec{x}=[\vec{p}_{I},\vec{q}_{IB},\vec{v}_{I}]^{T}$, where $\vec{p}_{I}\in\mathbb{R}^{3}$ is the position, $\vec{q}_{IB}\in\mathbb{SO}(3)$ is the unit quaternion that describes the rotation from the body to the inertial frame, and $\vec{v}_{I}\in\mathbb{R}^{3}$ is the linear velocity vector. $\vec{\omega}_{B}\in\mathbb{R}^{3}$ are the bodyrates in the body frame $B$. For ease of readability, we drop the frame indices, as they remain consistent throughout the description. The nominal dynamic equations are given by:
[…]
@@ IV Simulation Experiments
We implemented DreamerV3 in PyTorch, leveraging the dreamerv3-torch [^1] https://github.com/NM512/dreamerv3-torch open-source implementation and building on top of the stable-baselines3 library [47]. Our experiments were conducted within a high-fidelity simulation environment combining Flightmare [48] and Agilicious [40] for realistic quadrotor dynamics and track generation. To enable fast, real-time rendering and direct access to the image feed during training, we integrated the Habitat simulator [49, 50], enabling training with the renderer in the loop at several thousand frames per second. This same training environment has been successfully employed in prior research [13].
Our method’s performance is benchmarked against a model-free Proximal Policy Optimization (PPO) baseline [51, 4, 3, 12]. The PPO baseline used four-layer Multi-Layer Perceptrons (MLPs) with 768 neurons per layer for both the actor and critic networks.
The simulated quadrotor’s physical parameters were consistent across all experiments: mass $m=0.6$ kg, diagonal inertia matrix $J=\text{diag}([0.002410,0.001800,0.003759])$ kg m2, rotor torque constant $\kappa=0.022$, and arm length $0.14$ m. The maximum rotor thrust was limited to $4.0$ N, resulting in a thrust-to-weight ratio of $2.7$.
@@ IV-A Results
We conducted training experiments on our agent using two distinct reinforcement learning algorithms: DreamerV3 and PPO. These experiments were carried out across three challenging drone racing tracks: Circle, Kidney, and Figure 8. The results of these experiments are visualized in Figure 3. The top row of the figure presents the overall reward evolution for each track, comparing the performance of both DreamerV3 and PPO. The bottom row provides a zoomed-in view of the PPO reward curves, for a more detailed examination of their evolution. Each run was conducted with 5 different random seeds. The resulting plots depict the average reward across these 5 runs, represented by a solid center line, and the shaded area surrounding the line indicates the standard deviation. Our findings indicate that DreamerV3, by leveraging a learned world model, effectively learns directly from pixel observations. This enables the agent to acquire policies that are capable of agilely flying through each track. In contrast, PPO, which relies solely on direct interaction with the environment, struggles to converge to high-reward solutions. Its policies often lead to policies that fail to execute any meaningful flight maneuvers, resulting in consistently low reward values. This aligns with the findings presented in [13], where the authors encountered similar challenges in training end-to-end PPO policies directly from pixel observations. To overcome these difficulties, they adopted a hybrid approach, combining imitation learning with subsequent reinforcement learning fine-tuning.
@@ IV-B Perception-aware emergent behaviour
Figure 5: Ablation study on the perception aware behaviour of our policies. Top: DreamerV3 policy trained on pixel observations in an environment where the only rendered gates are the actual gates. As indicated by the black arrows (representing camera direction), the platform predominantly focuses its attention on the next gate. Bottom: We introduce two additional gates to the rendering engine (marked in red color). These gates are not required to be traversed but serve as a valuable source of information for platform localization. Consequently, the policy’s behavior shifts, and the platform now distributes its attention more evenly across both the actual and the extra gates.
An additional interesting observation is that after training, the agent exhibits a consistent tendency to orient the camera towards the gates, a behavior that naturally emerged during training rather than being explicitly incentivized by the reward function, as it is generally done in previous works [12, 13]. This emergent behavior can be attributed to two causes. First, the fact that we are optimizing end-to-end from pixels to commands, therefore allowing for closing the action-perception loop. And second, to the inherent characteristics of the visual observations. As depicted in Figure 4, the gates remain visually clear throughout the entire track, while the background details become increasingly blurred due to the downsampling of the input image to $64\times 64$ pixels. This suggests that the policy strategically prioritizes focusing on the information-rich gates, which are essential for successful navigation. To reinforce this hypothesis, we have conducted an additional ablation experiment where we place two additional gates in the periphery of the Figure 8 track. These gates are there only in the rendering, but do not need to be passed through. Fig. 5 shows the same policy trained with and without these additional gates. In the top part of Fig. 5, we show the policy trained without the extra gates. As indicated by the black arrows (representing camera view direction), the platform predominantly focuses its attention on the following gates. In the bottom part of the figure, we show the policy trained with the two additional gates to the rendering engine (marked in red color). As it can be seen, the policy’s behavior changes, and the platform now distributes its attention more evenly across both the actual and the extra gates.
@@ V Real-world Experiments
Figure 6: Real-world experimental setup. Our drone is equipped with a RF receiver similar to that used by professional human pilots. The observations are onboard RGB images, similar to those seen by a human pilot.
V-A Hardware Setup
We deploy our approach in the real world using a high-performance racing drone with a maximum thrust-to-weight ratio (TWR) of 5.78. However, for our experiments, we have limited the TWR to 2.7. We use a modification of the Agilicious platform [40] for the real-world deployment. We have replaced the onboard computer with an RF receiver, which is connected directly to the flight controller [^2] https://www.betaflight.com and takes care of parsing the collective thrust and bodyrate commands from the offboard computer. This configuration is similar to the one used by professional drone racing pilots. This quadrotor and racing setup is shown in Fig. 6. For the deployment in the real world, we use a hardware-in-the-loop (HIL) setup, where the images are obtained from the habitat simulator and fed into the network, and the commands produced are sent directly to the real drone platform. This way, we have the real world dynamics in the loop, which allows us to asses the sim-to-real gap and our policy performance when tested in the real system.
@@ V-B Results
We demonstrate the efficacy of our policies by deploying them in the real world on the Figure 8 track. Figure 7 presents a comparative analysis of the simulated and real-world trajectories and speed profiles for this track. A strong similarity is observed between the two, indicating a minimal sim-to-real gap in our dynamic model. Moreover, the camera axis, visualized by black arrows in Figure 7, aligns closely between the simulated and real-world scenarios. As already mentioned in Section IV-B, both the simulated and deployed policies show the same perception-aware behavior: they keep the camera view aligned with the next gate, even if there are no reward terms guiding or incentivizing it. The real world deployment is also shown in the supplementary video.
@@
VI Limitations
The proposed model-based RL approach has shown great success at learning complex behaviours from pixels to commands at agile flight, also demonstrating successful real-world deployment. However, there are some limitations, mainly related to the fact that our approach needs a rather large amount of computation for both training and deployment. To be able to train a policy that performs to convergence (for 20 million steps, as shown in Fig. 3, our architecture takes an average time of 240 hours. This time comes primarily from the forward and backward propagation through the RSSM world model, and not from the rendering. This means that more efficient implementations of DreamerV3 hold the potential to reduce the training times. For example, using the JAX implementation [^3] https://github.com/danijar/dreamerv3 instead of the PyTorch one can potentially reduce the training times by a factor of 3-4 times. Another key limitation of our current approach relies in the substantial computational resources for real-time operation. Therefore for real-time deployment our approach necessitates offboard processing, which is a limitation, but can also be seen as an enabler for future research, which allows us to leverage the rapid advancements in high-performance computing, including access to cutting-edge foundation models and powerful cloud infrastructure, without being constrained by the limitations of onboard processing capabilities.
VII Conclusion
This paper introduces a novel application of MBRL using the DreamerV3 architecture to address the challenge of vision-based agile quadrotor flight. Our approach eliminates the need for intermediate representations, explicit state estimation, or extensive bootstrapping by mapping raw RGB pixel inputs directly to control commands.
Our results demonstrate the efficacy of DreamerV3 by executing a drone racing task, both in simulation and the real world. Compared to model-free RL methods (in particular PPO), our approach achieves significantly higher sample efficiency and produces policies capable of navigating challenging race tracks at high speeds (up to 8 m/s). Additionally, the emergent perception-aware behavior observed in trained policies highlights the potential of end-to-end learning systems (without intermediate representations) to not only allow perception to optimize action, but also enable action to optimize perception. In general, our work provides a step forward applying pixel-to-command model-based RL to real-world mobile robotics.
Acknowledgments
This work was supported by the European Union’s Horizon Europe Research and Innovation Programme under grant agreement No. 101120732 (AUTOASSESS) and the European Research Council (ERC) under grant agreement No. 864042 (AGILEFLIGHT).

### v2

@@ III-D Model-based Reinforcement Learning: DreamerV3
III-D1 World Model training
@@
  • Recurrent Sequence Model: A recurrent sequence model, parameterized by the recurrent state $\bm{h}_{k}$, predicts the evolution of the latent representation. Given the previous latent state $h_{k-1},z_{k-1}]$ and the action $\bm{a}_{k-1}$, it predicts the current recurrent state $\bm{h}_{k}$.
@@
III-D2 Actor-Critic Training
@@
Figure 3: Reward evolution by number of steps for three different tracks: Circle track, Kidney Track and Figure 8 track. The training performance of DreamerV3 is shown in blue, for PPO in orange and for SAC in red. We show that none PPO nor SAC are able to achieve any considerable training in 20 million environment interactions, while DreamerV3 is able to train to convergence for the three tracks.
@@ IV Simulation Experiments
We implemented DreamerV3 in PyTorch, leveraging the dreamerv3-torch open-source implementation and building on top of the stable-baselines3 library [27]. Our experiments were conducted within a high-fidelity simulation environment combining Flightmare [32] and Agilicious [6] for realistic quadrotor dynamics and track generation. To enable fast, real-time rendering and direct access to the image feed during training, we integrated the Habitat simulator [29, 25], enabling training with the renderer in the loop at several thousand frames per second. This same training environment has been successfully employed in prior research [39].
Our method’s performance is benchmarked against two model-free baselines: Proximal Policy Optimization (PPO) [34, 33, 15, 8] and Soft Actor Critic (SAC) [9]. Both the PPO and SAC baselines used a CNN followed by four-layer Multi-Layer Perceptrons (MLPs) with 768 neurons per layer for both the actor and critic networks. The physical parameters of the quadrotor used were consistent across all experiments: mass $m=0.6$ kg, diagonal inertia matrix $J=\text{diag}([0.002410,0.001800,0.003759])$ kg m2, rotor torque constant $\kappa=0.022$, and arm length $0.14$ m. The maximum rotor thrust was limited to $4.0$ N, resulting in a thrust-to-weight ratio of $2.7$. The dynamics, platform and parameters are the same one as in  [8].
@@ IV-A Results
We conducted training experiments on our agent using three distinct reinforcement learning algorithms: DreamerV3, PPO and SAC. These experiments were carried out across three challenging drone racing tracks: Circle, Kidney, and Figure 8. The results of these experiments are visualized in Figure 3. The top row of the figure presents the overall reward evolution for each track, comparing the performance of both DreamerV3, PPO and SAC. Each run was conducted with 5 different random seeds. The resulting plots depict the average reward across these 5 runs, represented by a solid center line, and the shaded area surrounding the line indicates the standard deviation. Our findings indicate that DreamerV3, by leveraging a learned world model, effectively learns directly from pixel observations. This enables the agent to acquire policies that are capable of racing through each track. In contrast, PPO, and similarly SAC, which rely solely on direct interaction with the environment, struggle to converge to high-reward solutions. Furthermore, the model-free policies fail to execute any meaningful flight maneuvers, resulting in consistently low reward values. An important observation is that our system is trained using curriculum learning, where the parameter $b_{2}$ in Eq. (1) is initially set to $0.0$ and gradually increases once the total reward surpasses $50.0$. This explains the oscillatory reward values observed in Fig. 3, as the system adjusts to the changing rewards.
Our experimental results align with the findings presented in [39], where the authors encountered similar challenges in training end-to-end PPO policies directly from pixel observations. To overcome these difficulties, they adopted a hybrid approach, combining imitation learning with subsequent reinforcement learning fine-tuning. In contrast, using a model-based approach, we are able to completely avoid such additional efforts. We note, however, that this comes at the cost of a significantly longer training time than the baselines, reaching 240 hours.
@@ IV-B Perception-aware emergent behaviour
Figure 5: Ablation study on the perception aware behaviour of our policies. Top: DreamerV3 policy trained on pixel observations in an environment where the only rendered gates are the actual gates. As indicated by the black arrows (representing camera direction), the platform predominantly focuses on the next gate. Bottom: We introduce two additional gates to the rendering engine (marked in red color). These gates are not required to be traversed, and are not part of the objective, but serve as a valuable source of information for platform localization. Consequently, the policy’s behavior shifts, and the platform now distributes its camera view more evenly across both the actual and the extra gates.
An additional interesting observation is that after training, the agent exhibits a consistent tendency to orient the camera towards the gates, a behavior that naturally emerged during training rather than being explicitly incentivized by the reward function, as it is generally done in previous works [8, 39]. This emergent behavior can be attributed to the fact that we are optimizing end-to-end from pixels to commands, therefore allowing for closing the action-perception loop. As depicted in Figure 4, the gates remain visually clear throughout the entire track, while the background details become increasingly blurred due to the downsampling of the input image to $64\times 64$ pixels. This suggests that the policy strategically prioritizes focusing on the information-rich gates, which are essential for successful navigation. To reinforce this hypothesis, we have conducted an additional ablation experiment where we place two additional gates in the periphery of the Figure 8 track. These gates are there only in the rendering, but do not need to be passed through. Fig. 5 shows the same policy trained with and without these additional gates. In the top part of Fig. 5, we show the policy trained without the extra gates. As indicated by the black arrows (representing camera view direction), the platform predominantly focuses on the following gates. In the bottom part of the figure, we show the policy trained with the two additional gates to the rendering engine (marked in red color). As it can be seen, the policy’s behavior changes, and the platform now distributes its camera view more evenly across both the actual and the extra gates. In the supplementary video we show the first person rendered view of the two policies depicted in Fig. 5 deployed in a simulated environment.
@@ V Real-world Experiments
V-A Setup
The software setup is identical to the one used in the simulation experiments, explained in Section IV. Regarding the hardware, we use a modification of the Agilicious platform [6] for the real-world deployment. We have replaced the onboard computer with an RF receiver, which is connected directly to the Betaflight flight controller and takes care of parsing the collective thrust and bodyrate commands from the offboard computer. Our hardware setup is the same setup as in [8], similar to the one used by professional drone racing pilots. For the deployment in the real world, we use a hardware-in-the-loop (HIL) setup, where the images are rendered using the habitat simulator and fed into the network, and the commands produced are sent directly to the real drone platform. This way, we have the real world dynamics in the loop, which allows us to asses the sim-to-real gap and our policy performance when tested on the real system.
@@ V-B Results
We demonstrate the efficacy of our policies by deploying them in the real world on the Figure 8 track. This experiment is shown in the supplementary material video at https://www.youtube.com/watch?v=nctQ2rxZnIc. Figure 6 presents a comparative analysis of the simulated and real-world trajectories and speed profiles for this track. A strong similarity is observed between the two, indicating a minimal sim-to-real gap in our dynamic model. Moreover, the camera axis, visualized by black arrows in Figure 6, aligns closely between the simulated and real-world scenarios. As already mentioned in Section IV-B, both the simulated and deployed policies show the same perception-aware behavior: they keep the camera view aligned with the next gate, even if there are no reward terms guiding or incentivizing it. To the best of our knowledge, our HIL demonstration marks the first RL approach to learn drone racing directly from pixel inputs to control commands, effectively closing the loop between perception and action. The supplementary video shows the deployment of our system in the real world.
@@
VI Conclusion
This paper presented a MBRL approach using DreamerV3 to train end-to-end visuomotor policies for agile drone flight. Our method learns directly from raw pixel inputs, removing the need for intermediate representations or imitation learning bootstrapping. Our experiments demonstrated that this approach is more sample-efficient than model-free baselines like PPO and SAC and results in emergent perception-aware behaviors without explicit reward shaping.
We validated the learned policies in simulation and on a physical quadrotor using a hardware-in-the-loop (HIL) setup. While this confirms the policy’s effectiveness on the real quadrotor dynamics, additional challenges remain to transfer to real camera pixel observations. Additionally, the training process is computationally intensive, requiring approximately 240 hours to converge. Future work should focus on using photorealistic simulators or exploring other observation modalities that allow for deployment directly real pixels, and on improving the computational efficiency of the world model. Overall, our work is a promising step towards applying pixel-to-command MBRL on real-world robotic systems.